Hi i really need your help with this equation.I really would appreciate it!I need help with figuring what "x" is in the equation.The equation is 1/4x-6=15.Please could you show me this step by step?Thank you!!!

Found 2 solutions by checkley75, ssr1292:
Answer by checkley75(3666)

1/4x-6=15
1/4x=15+6
1/4x=21 now cross multiply
x=84 answer.

Answer by ssr1292(1)

1/4x-6=15
+6=+6
1/4x=21
4/1*1/4x=21*4/1
x=84